OpenStudy (anonymous):

If the money and interest are left in the bank, how much is the total savings after three years for $2,000 at 11% annually compounded interest? A. $348.18 B. $1,390.35 C. $2,735.26 D. $3,200.50

OpenStudy (amistre64):

3 years is a pretty short time, so a long version wont be that difficult to work thru

OpenStudy (amistre64):

Y1 = B + Br Y2 = (B + Br)+(B + Br)r Y3 = ((B + Br)+(B + Br)r) + ((B + Br)+(B + Br)r)r

OpenStudy (amistre64):

as a rule of thumb, we shouldnt be less than what we started with ... so A. $348.18 B. $1,390.35 are just pointless

OpenStudy (anonymous):

C?

OpenStudy (amistre64):

C would seem like a safe bet, but i havent tried to calculate it Y3 simplifies to: B(1+r)^3 so 2000(1.11)^3

OpenStudy (amistre64):

yep